OpenStudy (anonymous):

(b/b^2+6b+50)+(5/b^2+6b+5)

OpenStudy (anonymous):

add

OpenStudy (anonymous):

\[\left(\begin{matrix}b \\ b^2+6b+5\end{matrix}\right)+\left(\begin{matrix}5 \\ b^2+6b+5\end{matrix}\right)\]